Calculate the uncertainty in wavelength of a light emitted when an atom undergoes transition from an excited state that lasts for 10-3 second.



1
Expert's answer
2022-05-02T13:43:23-0400

"Et=\frac h2,"

"E=\frac h{2t},"

"E=5.3\cdot 10^{-32}~J."